the probability that a college student owns a microwave is 0.63 , that a college student owns a refrigerator is 0.25 and that a college student owns both is 0.16 find the probability that a college students owns a microwave or refrigerator
Answer by ikleyn(52781)

.

P(microwave OR refrigerator) = P(microwave) + P(refrigerator) - P(microwave AND refrigerator) = 0.63 + 0.25 - 0.16 = 0.72.    ANSWER